LOFAR Newsletters April 2015

Distributed on 20 April 2015.

 

 

Array status:

----------------------

- 38 stations operational in the Netherlands: 24 core stations and 14 remote stations


- 9 international stations operational: DE601, DE602, DE603, DE604, DE605, FR606, SE607, UK608, DE609


- 3 new international stations are being built in Poland and should become part of the array by the end of 2015


- It was discovered that the rubidium clock was not properly installed at DE609 – it has been fixed now. New calibration tables produced. Similar problems might be affecting other international stations – investigations are in progress.


- RFI sources nearby DE609 affect the quality of the LBA data of this station. Investigation is in progress to solve this issue.


- The overview of non-operational antenna elements for LBA and HBA is available here. The maintenance season is starting. A few international stations will be visited during the coming few weeks.


- Station calibration: new mode 7 calibration tables (HBA HIGH) have been updated for all stations in Feb 2015. Mode 5 tables have been updated for the Remote Stations – the rest of the array will follow next week. Mode 1-3 (LBA) calibration runs are in progress. The current station calibration status can be found here.

  

Observing System Status:

------------------------------------ 

- The observing system was mostly stable during the period Feb-Apr 2015. However, a few failures at COBALT nodes, as well as the instability of processing and ingest systems caused delays in the exploitation of the Cycle 3 observing program and in digesting data out of CEP2.


- During a software roll out on January 26th 2015, the information on broken tiles and antenna elements, which is stored in a LOFAR internal database, was inadvertently erased. Consequently, from that moment onward the list of malfunctioning antenna elements stored in the LOFAR Measurement Sets (MS's) only contained newly diagnosed malfunctioning elements. This problem was solved on February 25th 2015. Since then, the MS's contain again the proper broken tile and element information. The PI’s of all affected projects have been notified. A script to correct the data has been provided to them. Full description of the issue is given here

 

- On the LTA side, we have started to utilize the Target storage infrastructure in Groningen as an LTA data facility. Currently we are migrating data to the production area. Efforts are underway to allow LOFAR science project teams to ingest their data into the LTA.

 

Software Development Status:

------------------------------------------

Message Queues, based on the QPID framework, have been introduced in the LOFAR software environment. The first deployment includes a solution to a known limitation in the handling of information feedback from observations and processing pipelines. In the coming months we will look at applying Message Queues in other areas to help improve performance and reliability of inter process communication, an essential improvement to the LOFAR system to allow further implementation of responsive services. Other activities include a fix for a 0.5 second integration issue for COBALT and improvements & fixes in pulsar, selfcal, and long baseline pipelines.

 

CITT Status:

-----------------

The CITT has booked progress creating a pipeline version of the facet calibration technique. This pipeline will be tested during the facet calibration workshop in Leiden. Integration of the new AWimager is on track, thanks to the release of a new casacore. We aim for a release of the new imager in LofIm at end of June, while it is already available on the CITT branch.

 

 MSSS status:

------------------

The Multifrequency Snapshot Sky Survey (MSSS) is now working toward its first HBA data release. To coordinate the first batch of science publications, a workshop is being organized in Dwingeloo on 4-5 June 2015 (immediately following the LOFAR Users and Science Meetings). Those interested in participating can register at the workshop webpage by 1 May 2015. Work on the MSSS-HBA catalog is currently focused on verification of the flux scale, as has been described at a series of recent LOFAR Status Meetings. This flux scale investigation is also being used to provide an improved beam model to the LOFAR community.

 

 

Observing Programmes and miscellaneous items:

--------------------------------------------------------------------

- Cycle 2 observing programme: 90% complete. Remaining 10% is being observed with second priority during the current Cycle. Note: unfinished projects will not continue after May 14th 2015.


- Cycle 3 observing programme: ~ 75% complete. Observing schedule advertised here. All Cycle PI's are solicited to check the schedule and get in contact with Science Support at sciencesupport [at] astron [dot] nl in case of questions/remarks. Note: it will not be possible to complete the Cycle 3 observing program by May 14 2015. Unfinished projects will continue during Cycle 4, with second priority.


-  The Cycle 4 proposal submission deadline was on March 11 2015.  38 documents were received, of which 37 are regular proposals and 1 is an envelope sheet.  The oversubscription rates are 1.4 for observing and 1.7 for processing. The review process will finish with the PC meeting, planned for May 6-7 2015, at ASTRON.


- As of March 1 2015, all LOFAR data in the LTA that have passed the proprietary period(s) have become public. Users can submit requests to download such data. The table of contents of the LTA is fully public at any time. Information on how to use the LTA can be found here.  The LTA is accessible here.  

 

 

CEP news:

----------------

- CEP2:

 

a) The RAID controller of locus084 crashed on March 13 2015 and all data on this node got lost.

b) Disk space situation is often critical, especially in case of instability of the processing and ingest systems and after week-ends.

 

- CEP3:

 

a) lof011 experienced a corrupted file system during the stop day on 7 April 2015. All data on this node got lost.

b) Several requests for extending the access to CEP3 are submitted by users. However, nodes seem to be under-used or not used at all during the originally awarded period.  Monitoring of node usage is active now and these statistics will be used to evaluate future requests. 

 

Calendar next LOFAR activities:

-------------------------------------------------------------------

- A reconfiguration of the LOFAR network is planned from May 18 till May 29 2015. During the network reconfiguration, access to/functionality of any system in the LOFAR domain is not guaranteed. Note: access to the CEP systems (CEP2 and CEP3) will not be possible during the reconfiguration period


- Next LSM's : 29/04, 13/05, 27/05, 10/06, 24/06, 08/07


- Next Stop days: 02/06, 07/07


- Next software roll outs: 29/06, 31/08


- LOFAR Network Reconfiguration: 18-29/05


- 2nd LOFAR Users Meeting + 2015 LOFAR Science Workshop: June 1-3 2015, Assen (NL):

a) Meetings page here

b) Registration and abstract submission deadline: May 1 2015

 

- LOFAR MSSS Science Workshop 2015: June 4-5, ASTRON

a) Meetings page here

b) Registration and abstract submission deadline: May 1 2015


- Next LOFAR bulletin: June 2015

 

Design: Kuenst.    Development: Dripl.    © 2020 ASTRON